OpenStudy (anonymous):

Explain how to solve the equation 3t + 11 = –10. Show that your solution works. (4 points)

OpenStudy (anonymous):

3t = -10 -11 3t = -21 t = -21/3 t = -7

OpenStudy (anonymous):

SS = { -7}

OpenStudy (anonymous):

Thanks you!!

OpenStudy (anonymous):

Gimme a Medal girl :)